OpenStudy (anonymous):

Find the volume of the cylinder in terms of Pi. h = 6 and r = 3

OpenStudy (lgbasallote):

volume = pi r^2 h volume = pi (3)^2 x 6 can you do the rest?

OpenStudy (lgbasallote):

it says in terms of pi so you dont need to convert pi..just put it as \(\pi\) okay?

OpenStudy (anonymous):

169.6?

OpenStudy (lgbasallote):

no no...like i said DONT convert pi because it says in terms of pi

OpenStudy (lgbasallote):

so just square 3..then multiply that to 6...then just attach a \(\pi\) symbol then you're done

OpenStudy (anonymous):

54!

OpenStudy (lgbasallote):

remember i said attach \(\pi\)? therefore it's \(54\pi\)